OpenStudy (anonymous):

The sum of two numbers is 47. The greater number is 9 more than the smaller number. Which equation can be used to solve for the smaller number?

OpenStudy (anonymous):

x+9=47

OpenStudy (anonymous):

can someone please control the freak that is innapropriate please!

OpenStudy (mathstudent55):

The smaller number is x. The larger number is x + 9. Now add the two expressions above and set the sum equal to 47.

OpenStudy (anonymous):

these are the equations x + (x + 9 )= 47 x - ( x + 9 n)= 47 x(x + 9) = 47 x(x - 9) = 47

OpenStudy (anonymous):

im stumped. i have no idea. sry